### v1

# Controlling dynamical systems into unseen target states using machine learning

Abstract
We present a novel, model-free, and data-driven methodology for controlling complex dynamical systems into previously unseen target states, including those with significantly different and complex dynamics. Leveraging a parameter-aware realization of next-generation reservoir computing, our approach accurately predicts system behavior in unobserved parameter regimes, enabling control over transitions to arbitrary target states. Crucially, this includes states with dynamics that differ fundamentally from known regimes, such as shifts from periodic to intermittent or chaotic behavior. The method’s parameter-awareness facilitates non-stationary control, ensuring smooth transitions between states. By extending the applicability of machine learning-based control mechanisms to previously inaccessible target dynamics, this methodology opens the door to transformative new applications while maintaining exceptional efficiency. Our results highlight reservoir computing as a powerful alternative to traditional methods for dynamic system control.
Introduction
Controlling dynamical systems is a fundamental challenge in engineering and science, crucial for maintaining stability, optimizing performance, and enabling adaptability across a broad spectrum of applications. From aerospace engineering, where precise control of aircraft and spacecraft is vital [1, 2], to industrial automation, where maintaining efficiency and safety in manufacturing processes is essential [3], the ability to effectively control dynamical systems is driving significant advancements among different fields [4, 5, 6, 7, 8]. In biomedical engineering, controlling physiological systems and biomedical devices improved patient care[9, 10, 11]. In the realm of energy systems, it enabled the efficient operation of power grids and the integration and optimization of renewable energy sources [12, 13]. However, the increasing integration of distributed renewable energy sources poses challenges to the robustness of power grids. Consequently, it is imperative to develop new control strategies tailored for future micro grids to ensure stability and reliability [14]. Traditionally, control methods such as Proportional-Integral-Derivative (PID) controllers, state-space control, and model predictive control (MPC) have been employed to manage these systems. While effective in many cases, these approaches often rely on accurate mathematical models and can struggle with the complexities inherent in nonlinear, high-dimensional, and time-varying systems [15, 16, 17]. In recent years, the advent of machine learning introduced powerful new tools for predicting and controlling dynamical systems. Techniques such as deep learning, neural ordinary differential equations and reinforcement learning showed promise in modeling complex systems and developing control strategies [18, 19, 20]. However, these methods often require extensive computational resources and can be difficult to interpret and implement. Reservoir computing (RC) emerges as a promising candidate in this landscape, offering a simpler yet effective approach for handling the temporal dependencies and nonlinearities of dynamical systems. Reservoir computing leverages a fixed, high-dimensional dynamical system, the reservoir, to process temporal data. The key advantage of reservoir computing lies in its separation of dynamic processing, managed by the reservoir, from the training process, which optimizes only the readout layer. This significantly reduces the computing costs and makes it orders of magnitude more efficient than comparable algorithms[21, 22]. Recent advancements in reservoir computing led to the development of various reservoir architectures, enhancing their predictive capabilities and computational efficiency. Traditional reservoir computing employs a randomly generated network at its core [23], providing robust performance in learning and predicting temporal patterns. Minimal reservoir computing simplifies this approach significantly by using a block-diagonal matrix structure, which lowers computational complexity and minimizes training data requirements [24]. Additionally, the next-generation reservoir (NGRC) computing architecture, originating from nonlinear vector autoregression, utilizes a deterministic library of unique monomials derived from the input data and time shifts of it [25]. Studies comparing these architectures demonstrated their effectiveness across various applications. For instance, the prediction performance of these architectures on chaotic systems showed that minimal reservoir computing and next-generation reservoir computing outperforms traditional methods in scenarios with limited data availability, providing more accurate and reliable predictions [24]. Further research showed that next-generation reservoir computing can achieve comparable predictive performance to the already data-efficient traditional reservoir computing architectures while requiring only a fraction of the training data, about $10$ to $100$ times less [25, 26, 27]. These innovations make reservoir computing as a framework more flexible and effective, especially in scenarios with limited data and strong nonlinearity, opening up its use in various fields and problems. One field is the control of dynamical systems. In [28], traditional reservoir computing is used to control dynamical systems into arbitrary periodic, intermittent or chaotic target states without requiring knowledge of the underlying governing equations. To achieve successful control, the reservoir computer needs to accurately learn the statistical behavior of the target dynamic. Then, a control force based on the RC prediction and the current system state can be applied on the system, forcing it into the desired target state of the RC. In [27] the control performance of this approach was compared to that with NGRC at its core. While both methods demonstrate similar performance in reproducing the climate of the target states during control, NGRC outperformed RC by requiring approximately ten times less training data to achieve these results. Here we leverage this method to control dynamical systems into previously unseen target states by extrapolating these states from few training data samples using parameter-aware NGRC[29]. Further, the parameter-aware approach allows for non-stationary prediction. With that we introduce a novel non-stationary controlling method that ensure smooth transitioning from the current system state to the desired target state.
Results
Prediction of unseen parameter states
We assess the performance of the proposed method on the Lorenz system by varying the bifurcation parameter $\rho$ within the range previously analyzed for the prediction of unseen states using traditional reservoir computing [30]. For $\rho>100$, the system exhibits periodic behavior. As $\rho$ decreases below 100, it undergoes period-doubling bifurcations and transitions to chaotic dynamics. The bifurcation behavior is visualized in red within Fig. 1 $\mathbf{A}$. We use the parameter-aware NGRC approach, proposed in [29], to learn the Lorenz system using four periodic and stationary training data samples with parameters $\rho_{\text{train}}=100.0,100.1,100.2$ and $100.3$. After successful training we are able to extrapolate and interpolate the system behavior in unseen parameter regions of $\rho$, showing in Fig. 1 $\mathbf{B}$. For each parameter, we use a total data length of $5.000$ time steps for training the NGRC architecture. Compared to the $250.000$ training data points and additional $50.000$ synchronization data points per parameter used in [30], it becomes significant that the NGRC approach is highly data efficient when learning such dynamics.
Control into unseen parameter states
Accurate prediction of the dynamics of previously unseen states now allows for the application of the control mechanism described in [28, 27] to control the system dynamics toward these states. A control force is applied, defined as the product of a force scaling parameter $K$ and the difference between the system’s current state at $\rho_{\text{test}}$ and the predicted state of the target dynamics at $\rho_{\text{pred}}$. This control force directs the system’s dynamics at each time step to converge with the predicted dynamics. For example, the dynamic at $\rho_{\text{test}}=100.7$ can be controlled to align with any $\rho$ value within the analyzed bifurcation parameter range, as shown in Fig. 1 $\mathbf{C}$. This control mechanism is not limited to NGRC and is applicable to any method capable of accurately extrapolating and interpolating system dynamics.
Figure 1: Controlling dynamics into unseen states. A) Segments of the Lorenz bifurcation diagram (red). Four training samples and their bifurcation parameters are used for training (light green). B) Predicted bifurcation diagram for unseen parameter regions using the trained model. C) Controlled bifurcation diagram demonstrates the ability of the control mechanism to control system states into arbitrary and unseen target states using the models predictions.
Figure 2: Non-stationary transition duration influences system behavior. The orange line in (a) and (b) shows the parameter transition over time, while the red line represents the corresponding system dynamics from simulations. In (a), an abrupt parameter transition causes turbulent behavior after the transition, as also indicated by the increased running correlation dimension in the lower plot. In (b), a slow parameter transition leads to a smooth progression towards the target dynamics.
Non-stationary control
The parameter-aware NGRC approach has been shown to predict non-stationary time series effectively [29]. This enables not only direct control of the trajectory towards the target system but also introduces non-stationary transitions by controlling the system from $\rho_{\text{test}}$ through intermediate $\rho$ values. Figure 2(a) demonstrates that abrupt changes in $\rho$ (e.g. $100.5$ to $99.5$) can result in turbulent [31] behavior post-transition, while Figure 2(b) shows smooth transitions to target dynamics with gradual parameter changes. An examination of the occurrence of turbulent behavior during different transition times is conducted, considering the potential adverse effects of turbulence in practical and technological applications. We compare the non-stationary simulation behavior with the non-stationary predictions of the parameter-aware NGRC model and develop a control mechanism that completely eliminates the occurrence of turbulent behavior, regardless of the transition time. We analyze how the system simulations as well as the NGRC predictions behave when they transition from a periodic orbit at $\rho_{\text{start}}=100.5$ to a chaotic regime at $\rho_{\text{target}}=99.5$. Here, the transitions are performed with linearly changing $\rho$ values. Depending on the transition time and the location of the trajectory when the parameter transition starts, turbulent behavior might occur. We utilize a running correlation dimension[32, 33] (rcd) approach as an operational measure to detect changes in the system behavior. As shown in Figure 2(a), a global maximum of the $rcd$ forms during the transition phase, while in Figure 2(b), the global maximum of the $rcd$ is defined by the target dynamics at $\rho_{\text{target}}$. Since the $rcd$ of the target dynamics at $\rho_{\text{target}}$ does not exceed 1.21, we define $rcd_{\text{limit}}=1.21$ as the $rcd$-cut-off value. If the $rcd$ values of the non-stationary simulations or predictions remains below this threshold over time, we assume that the trajectory is smoothly transitioning into the target dynamics. Conversely, for simulations or predictions where the $rcd$ values exceed this threshold, we assume that they exhibit turbulent behavior, at least to some extent. Starting from 100 random initial conditions at $\rho_{\text{start}}$, we determine for different transition times how likely it is that the trajectories exhibit turbulent behavior during the transition phase to $\rho_{\text{target}}$. In Figure 3, we show that these probabilities are strongly dependent on the transition time and are approaching zero for long enough transition periods. The probabilities observed in the simulations are reproduced by the NGRC predictions, underscoring that the NGRC approach learns both stationary and non-stationary behavior patterns from only a few stationary data samples. Based on these findings a control mechanism is derived that avoids any possible turbulent behavior during the parameter transitions. As NGRC allows for efficient prediction of multiple realizations of test trajectories, we exploit that for each transition time at least some predictions are showing smooth transition behavior without turbulence occurring. These predictions can then be used to control the system dynamic’s at $\rho_{\text{start}}$, firstly for some time at $\rho_{\text{start}}$ and than for the smooth transition to $\rho_{\text{target}}$. By applying this approach to all 100 initialized systems at $\rho_{\text{start}}$ for each transition time, the control mechanism reduces the turbulence probability to zero. In Fig. 4 one of the controlled trajectories for a abrupt transition with no transition time is shown. To evaluate the quality of the controlled dynamics we compare the statistical climate of the actual system at $\rho_{\text{target}}$ to that of the $100$ controlled systems for each transition period. Therefore we measure the spatial complexity as well as the temporal complexity for each controlled system by calculating its correlation dimension[33] and its largest Lyapunov exponent[34], respectively. For the actual system and each controlled system, the measurements were carried out over $25.000$ time steps and are displayed in Fig. 5. The correlation dimension of the actual system is $cd=1.1378$, reproduced with an average correlation dimension of $cd=1.1382$ across all controlled systems. Similarly, the largest Lyapunov exponent of the actual system is $ly=0.0735$, while the average largest Lyapunov exponent across all controlled systems is $ly=0.0628$. Both metrics, particularly the correlation dimension, indicate that the statistical climate of the actual system is effectively predicted following each transition period, despite the model being trained on periodic trajectories and is extrapolating the emerging chaotic system behavior. With that the method enables smooth control into unseen target states avoiding adverse turbulent behavior during transitions.
Figure 3: Control with smooth transitions into unseen target states. Starting from $100$ initial conditions, the probability of turbulent behavior for the simulation (red) and the model prediction (green) is shown in red and green respectively, given different transition periods and a parameter transitions from a period orbit at $\rho_{start}=100.5$ to chaotic one at $\rho_{target}=99.5$. If predictions that show no turbulent behavior are taken as control dynamics, the turbulence probability can be reduced to zero.
Figure 4: Smooth control into target state . In (a) the initial periodic system dynamics at $\rho_{start}=100.5$ without control is plotted. In (b), the trajectory is smoothly controlled into the chaotic target dynamics at $\rho_{target}=99.5$ using the proposed control mechanism. A smooth transition is provided even if an abrupt transition to the target parameter is specified.
Figure 5: Statistical climate of controlled target states. Evaluation of the statistical climate of the $100$ controlled states at $\rho_{start}=99.5$ given each transition period shown in Fig. 3. In (a), the spatial complexity of the dynamics are measured using their correlation dimension. In (b), the temporal complexities are measured using the largest Lyapunov exponent.
Discussion
In this work, we proposed a model-free, data-driven machine learning methodology for smoothly controlling dynamical systems into previously unseen target states, including those with fundamentally different behaviors. By leveraging a parameter-aware machine learning model, we demonstrated the ability to interpolate and extrapolate system dynamics across unexplored parameter regimes. This approach enables transitions to target states that differ significantly from the observed dynamics, offering a novel framework for controlling complex systems. We utilize a stationary prediction of the target dynamic to define a threshold value, such a as the running correlation dimension, to quantify and avoid turbulent behaviors during transitions. By generating multiple non-stationary transition trajectories and filtering for predictions that don’t exhibit turbulence, we ensure smooth transition to the target dynamics. The control approach is generalizable under the condition that the underlying model extrapolates the bifurcation behavior correctly. The next-generation reservoir computing (NGRC) approach used stands out for its exceptional data efficiency, as further confirmed by our experiments on the Lorenz system. Compared to traditional reservoir computing, NGRC requires more than an order of magnitude less training data while maintaining accurate predictions[30]. This makes it particularly suitable for scenarios where data is sparse or expensive to collect. Additionally, NGRC’s concise architecture allows for hardware implementation on platforms such as FPGAs[35]. The current state in this field allows for NGRC hardware implementations with a single control target state. Notably, the method presented here extends the number of achievable target states per hardware implementation and enables non-stationary control, a key advantage for edge computing applications where real-time adaptability is crucial. However, several limitations warrant further discussion. A notable challenge of the underlying NGRC model is its susceptibility to stability issues during prediction if the hyperparameters are not well optimized. This sensitivity can pose challenges to its adaptability in more complex scenarios. Future work will focus on this issue by refining the two core building blocks of reservoir computing, the reservoir itself and its optimization algorithm, to further progress its adaptability while maintaining high data efficiency with architectural simplicity. Additionally, exploring the integration of NGRC with other machine learning paradigms and testing its performance across a wider range of more complex dynamical systems will further validate its applicability and uncover new potential use cases. By addressing these challenges, the proposed methodology holds promise as a transformative tool for controlling complex dynamical systems, opening pathways to applications where traditional methods fall short.
Methods
Next-Generation Reservoir Computing
Next-generation reservoir computing (NGRC) is a machine learning framework that enhances traditional reservoir computing by leveraging deterministic structures derived from nonlinear vector autoregression (NVAR). Instead of utilizing a randomly weighted internal network, NGRC constructs its feature vectors through a library of unique monomials of time-shifted input variables from the input data. This process involves concatenating $k$ past time series data points, each separated by $s$ time steps, to form a linear time-delay state vector. The final feature vector $\mathbf{r}_{i}$ is created by generating unique monomials of orders $O$ from this state vector, introducing nonlinearity into the model.[25] This way given a d-dimensional time series $\mathbf{X}$, a matrix of feature vectors $\mathbf{R}(\mathbf{X},k,s,O)=(\mathbf{r}_{1},\mathbf{r}_{2},\ldots,\mathbf{r}_{n})$ can be generated. Training is simplified by optimizing only the output layer $\mathbf{W}_{out}$ which maps the feature vectors onto the difference between the current point and the next point in time using ridge regression, such that
$\mathbf{x}_{i+1}=\mathbf{x}_{i}+\mathbf{W}_{out}\mathbf{r}_{i+1}.$ (1)
To parameterize the feature vector of the NGRC, the approach introduced in [29] is followed, wherein the parameter $\theta$ multiplied by a scaling factor $\gamma$ is added to each element of the feature vector, resulting in $\mathbf{r}_{i}^{\prime}=\mathbf{r}_{i}+\gamma\theta$. Additionally, $\mathbf{r}_{i}^{\prime}$ is expanded by including its own powers up to order $O_{\text{states}}$ to generate the parametrized feature vector $\widetilde{\mathbf{r}}_{i}$. To employ this architecture for multiple stationary dynamics during training, the concept of multifunctionality is leveraged [36]. Therefore, for each time series $\mathbf{X}_{m}$ and its parameter $\theta_{m}$, the parametrized feature vectors $\widetilde{\mathbf{R}}_{m}=(\widetilde{\mathbf{r}}_{m,1},\widetilde{\mathbf{r}}_{m,2},\ldots,\widetilde{\mathbf{r}}_{m,n})$ are generated along with their corresponding training targets $\mathbf{Y}_{m}$. To optimize the output layer with this setup, the parametrized feature matrices $\widetilde{\mathbf{R}}_{m}$ and the training target matrices $\mathbf{Y}_{m}$ are concatenated into single matrices $\widetilde{\mathbf{R}}_{M}$ and $\mathbf{Y}_{M}$, respectively. This allows the output layer $\mathbf{W}_{\text{out}}$ to be learned via ridge regression by solving the optimization
$\mathbf{W}_{\text{out}}=\mathbf{Y}_{M}\widetilde{\mathbf{R}}_{M}^{T}\left(\widetilde{\mathbf{R}}_{M}\widetilde{\mathbf{R}}_{M}^{T}+\beta\mathbf{I}\right)^{-1},$ (2)
where $\mathbf{I}$ is the identity matrix and $\beta$ is the regularization parameter. If the training is successful, the structure allows the parameter $\theta$ to change during the prediction, so that the prediction is performed under $\theta_{i}$
$\mathbf{x}_{i+1}=\mathbf{x}_{i}+\mathbf{W}_{\text{out}}\widetilde{\mathbf{r}}_{i+1}(\theta_{i+1}).$ (3)
In this way, both stationary and non-stationary dynamics can be predicted, even for parameter ranges that were not seen during training. The results were obtained using a NGRC architecture with parameters $k=2$, $s=47$, $O=[1,2,3]$, $O_{\text{states}}=3$, and $\gamma=0.18$. Additionally a bias feature of $1$ was included in the parameterized feature vector.
Control
Control theory refers to the study and application of methods for influencing the behavior of dynamical systems such that a desired target state is achieved. In the context of chaotic systems, these approaches are able to stabilize unstable periodic orbits by applying small perturbations through external forces. Traditional methods, such as OGY control [37] and delayed feedback control [38], typically rely on phase space techniques, requiring detailed knowledge of the system’s underlying equations or extensive datasets. These approaches are typically limited to steering systems toward simple dynamical target states. While various extensions have been proposed [39], including methods to "chaotify" periodic or synchronized dynamics [40], the ability to reach arbitrary dynamical target states has only recently been demonstrated [28]. In further research [27], this novel approach was used to compare traditional reservoir computing with next-generation reservoir computing (NGRC) in terms of the data requirements for effective control. NGRC demonstrated superior performance, requiring ten times less training data while achieving comparable control capabilities.
In this paper, we build on the control mechanism introduced in [28, 27] to guide dynamical systems into previously unseen target states. This general framework is not restricted to the next-generation reservoir computer employed here and can also be implemented with alternative methods, such as the minimal reservoir computer, traditional reservoir computing techniques, or other suitable machine learning techniques. The core idea is to train the machine learning method to predict the desired dynamics of the system, represented as the state $\textbf{X}(\rho_{i})$ at a given system parameter $\rho_{i}$. Notably, the dynamics of the desired state $\textbf{X}(\rho_{i})$ do not need to be recorded beforehand for training. Instead, the system can leverage the ability to predict unseen states, as explained in the previous Section, and be trained using the available data. If the system instead evolves to a different state Y — due to external influences or changes in its bifurcation parameters $\rho_{j}$ — an external force $\textbf{F}(t)$ is applied to guide the dynamics back to the desired state $\textbf{X}(\rho_{i})$. Without this intervention, the system would persist in state Y, following its natural trajectory $\textbf{u}(t)$.
To compute the appropriate control force $\textbf{F}(t)$, it is necessary to predict how the system trajectory $\textbf{u}(t)$ would evolve if it were in the desired state $\textbf{X}(\rho_{i})$. This hypothetical trajectory is defined as $\textbf{v}(t)$. Assuming the reservoir computer has been trained effectively, it can accurately forecast the future trajectory $\textbf{v}(t)$ as if the system were in state $\textbf{X}(\rho_{i})$. By comparing this predicted trajectory to the actual system evolution in state Y, the control force is computed based on the difference between the actual and hypothetical trajectories:
$\textbf{F}(t)=K(\textbf{u}(t)-\textbf{v}(t))\ ,$ (4)
where $K$ is a scaling parameter for the force which is set to $K=200$. This control force depends only on the observed actual system coordinates and the predicted coordinates, requiring no knowledge of the system’s underlying equations. This feature makes the approach highly versatile and applicable to real-world problems where precise mathematical models are unavailable. For a thorough analysis, we demonstrate this control strategy on a mathematical example system simulated using its known equations instead of observing a real system. In this case, the control force is incorporated by adding it to the system’s differential equations $\dot{f}$. The updated equations are then solved at each time step, incorporating the control force into the system’s evolution
[…]
To ensure a smooth transition to the system’s target state without turbulence occurring, a prediction of the target state must be selected that does not exhibit turbulent behavior. Multiple prediction realizations are generated for a given target state, and their correlation dimension is calculated. If the correlation dimension of a prediction is below the $rcd$-cutoff value, which represents the highest correlation dimension of the stationary target dynamics, it is considered free of turbulence and can be used for control. Depending on the reservoir computing method used, various strategies can generate multiple prediction realizations for a given target state. In our approach, we varied the initial conditions slightly and selected the turbulence-free prediction that starts closest to the system’s current coordinates when applying the control force.
Lorenz System
We selected the Lorenz system [41] as an example to demonstrate the functionality of the control mechanisms. In the analyzed bifurcation parameter range of $\rho$, the Lorenz system exhibits periodic behavior for $\rho>100$, undergoes period doubling for $\rho<100$ and transitions to chaotic behavior. The system of equations $\dot{f}(\mathbf{x}(t,\rho))$ is given by
$\dot{x}=\sigma(y-x),\\ \dot{y}=x(\rho-z)-y,\\ \dot{z}=xy-\beta z,$ (6)
where the parameters are set to $\sigma=10$ and $\beta=\frac{8}{3}$. These equations are solved for the simulation results using the fourth-order Runge-Kutta method with a time step size of $\Delta t=0.005$.
Running window correlation dimension
To determine whether the system undergoes a smooth transition during the non-stationary process, we employ a running correlation dimension approach computed over a defined window length. This technique captures the system’s dynamic spatial complexity over time. For periodic dynamics, the running correlation dimension values remain close to 1, whereas they exceed 1 when the dynamics exhibit chaotic behavior. In the case of the chaotic target dynamics of the example Lorenz system with $\rho=99.5$, this value remains below $1.21$. During non-stationary simulation, prediction, and control, we detect a non-smooth transition if the running correlation dimension surpasses $1.21$, indicating possible turbulence, while a smooth transition is considered if it remains below this threshold. For each windowed trajectory the correlation dimension $\nu$ quantifies the fractal dimensionality of the trajectory’s phase space [42].
$C(r)$ $={\frac{1}{W^{2}}\sum^{W}_{i,j=1}\theta(r-|\textbf{x}_{i}-\textbf{x}_{j}|)}$ (7)
where $\theta$ is the Heaviside function. The states are considered to be close to each other if their distance is less than the threshold distance $r$. The dimension $\nu$ is determined from the scaling relation:
$C(r)\propto r^{\nu}.$ (8)
We compute $\nu$ using the Grassberger–Procaccia algorithm [33] and use a window length of $W=500$ time steps for the Lorenz system.
Appendix
Acknowledgements
D. K. gratefully acknowledge the funding provided by Allianz Global Investors (AGI).
Author contributions statement
D. K. and A. H. developed the control algorithm and performed the experimental studies. C. R. initiated and supervised the work. All authors interpreted the findings and wrote the manuscript.